A pipe open only at one end has a fundamental frequency of 240 Hz. A second pipe, initially identical to the first pipe, is shortened by cutting off a portion of the open end. Now when both pipes vibrate at their fundamental frequencies, a beat frequency of 16 Hz is heard. How many centimeters were cut off the end of the second pipe? The speed of sound is 345 m/s.
Answer #1

fundamental frequency of pipe open at one end is given by:

f01 = V/(4L1)

Given that V = 345 m/sec and f0 = 240 Hz

So, L1 = V/(4*f01)

L1 = 345/(4*240) = 0.3594 m

Now for second pipe open at one end, beat frequency is 16 Hz, So

Since 2nd pipe is shortened, So it's fundamental frequency will increase, So

fundamental frequency of 2nd pipe will be:

Beat frequency = f02 - f01

f02 = 240 + 16 = 256 Hz

Now length of 2nd pipe will be:

L2 = V/(4*f02) = 345/(4*256)

L2 = 0.3369 m

So, change in length of pipe = dL = L1 - L2

dL = 0.3594 - 0.3369 = 0.0225 m

dL = 2.25 cm

Please Upvote.